OpenStudy (anonymous):

what is the Square root of X^13 please help

OpenStudy (anonymous):

\[\sqrt{x^{13}}\] two goes in to 13 6 times with a remainder of 1, so out comes \(x^6\)and one \(x\) stays in that gives \[\sqrt{x^{13}}=x^6\sqrt{x}\]

OpenStudy (anonymous):

Thanks a million! @satellite73